# Monocular 3D lane detection for Autonomous Driving: Recent Achievements, Challenges, and Outlooks

### v1

@@ Abstract
3D lane detection plays a crucial role in autonomous driving by extracting structural and traffic information from the road in 3D space to assist the self-driving car in rational, safe, and comfortable path planning and motion control. Due to the consideration of sensor costs and the advantages of visual data in color information, in practical applications, 3D lane detection based on monocular vision is one of the important research directions in the field of autonomous driving, which has attracted more and more attention in both industry and academia. Unfortunately, recent progress in visual perception seems insufficient to develop completely reliable 3D lane detection algorithms, which also hinders the development of vision-based fully autonomous self-driving cars, i.e., achieving level 5 autonomous driving, driving like human-controlled cars. This is one of the conclusions drawn from this review paper: there is still a lot of room for improvement and significant improvements are still needed in the 3D lane detection algorithm for autonomous driving cars using visual sensors. Motivated by this, this review defines, analyzes, and reviews the current achievements in the field of 3D lane detection research, and the vast majority of the current progress relies heavily on computationally complex deep learning models. In addition, this review covers the 3D lane detection pipeline, investigates the performance of state-of-the-art algorithms, analyzes the time complexity of cutting-edge modeling choices, and highlights the main achievements and limitations of current research efforts. The survey also includes a comprehensive discussion of available 3D lane detection datasets and the challenges that researchers have faced but have not yet resolved. Finally, our work outlines future research directions and welcomes researchers and practitioners to enter this exciting field.
@@ I INTRODUCTION
In recent advancements fueled by artificial intelligence, autonomous driving technology has seen rapid evolution, progressively reshaping the paradigms of human transportation. Equipped with an array of sensors, autonomous vehicles emulate human sensory capabilities such as vision and hearing to perceive their surroundings and interpret traffic scenarios for safe navigation[98]. Key among these sensors are LiDAR, high-resolution cameras, millimeter-wave radars, and ultrasonic radars[119, 55], which facilitate feature extraction and object classification combined with high-precision mapping to discern obstacles and the vehicular traffic landscape[114]. Visual sensors, the most extensively utilized in autonomous vehicles, serve as the primary means for environmental perception, encompassing lane detection, traffic light analysis, road sign detection and recognition, vehicle tracking, pedestrian detection, and short-term traffic prediction[11]. The processing and understanding of visual scenes in autonomous driving, which includes analysis of traffic lights, identification of traffic signs, lane detection, and the detection of nearby pedestrians and vehicles, underpin more robust and safer commands for maneuvers such as steering, overtaking, lane changes, or braking[82, 56]. This integration of sensory data and environmental understanding seamlessly transitions into the domain of scene understanding in autonomous driving, which is pivotal for advancing vehicular autonomy and ensuring road safety.
@@
Scene understanding represents one of the most challenging aspects within the domain of autonomous driving[57]. Without the capability for comprehensive scene understanding, navigating an autonomous vehicle safely through traffic lanes can be as daunting as navigating the world blindfolded for humans. Lane detection, in particular, stands out as a pivotal and challenging task within the realm of scene comprehension[105]. Lanes serve as the most common traffic elements on roads, acting as crucial markers that segment the roadways to ensure the safe and efficient passage of vehicles[79]. Lane detection technology, which automatically identifies road markings, is indispensable; autonomous vehicles lacking this capability could lead to traffic congestion and even severe collisions, thereby compromising passenger safety. Thus, lane detection plays a critical role in the ecosystem of autonomous driving. Unlike typical objects, lane markings occupy only a slender portion of the road scene and are widely distributed, making them uniquely challenging to detect[76]. The task is further complicated by a variety of lane markings, insufficient lighting, obstructions, and interference from similar textures, which are common in many driving scenarios, thus exacerbating the intrinsic challenges associated with lane detection.
@@
Lane detection methods based on monocular vision can be primarily categorized into traditional manual feature approaches and deep learning-based methods[58]. Early endeavors primarily focused on extracting low-level manual features[28, 41], such as edges and color information. However, these approaches often involve complex feature extraction and post-processing designs, and exhibit limited robustness in dynamically changing scenes. Traditional lane detection algorithms based on manual feature extraction commence by identifying characteristics such as color, texture, edges, orientation, and shape of the lane lines, followed by constructing detection models that approximate the lane markings as straight or higher-order curved lines[117]. Nevertheless, due to the lack of distinguishing features and poor adaptability to dynamic environments, traditional methods based on manual features are typically less reliable and more computationally expensive[96]. With the rapid evolution of deep learning, significant strides have been made in the fields of image classification, object detection, and semantic segmentation within computer vision, bringing innovative perspectives to the research of lane detection[97]. Deep neural networks (DNNs), rooted in deep learning, demonstrate profound capability in feature extraction from image data, with Convolutional Neural Networks (CNNs) being the most extensively applied. CNNs represent a specialized category of DNNs, characterized by multiple convolutional layers and bases, making them particularly suitable for processing structured data like visual images, and providing efficient feature extraction for various subsequent tasks[106]. In the context of lane detection, this translates to utilizing deep CNNs to extract high-level features in real-time frames, which are then processed by the model to accurately determine the positions of lane lines[31].
@@ I-A Background and Related Works
Due to advancements in deep learning technologies, researchers have developed numerous strategies that have significantly simplified, expedited, and enhanced the task of lane detection[78]. Concurrently, as deep learning becomes increasingly widespread and as new concepts continuously emerge[59], the domain of lane detection has seen a further specialization and refinement in its approaches. Reflecting on the mainstream research trajectories within this domain, camera-based lane detection methods can be principally segmented into two-dimensional (2D) and three-dimensional (3D) lane detection paradigms[12].
@@
2D lane detection. [35, 51] methods aim to accurately delineate the shape and position of lanes within images, primarily employing four distinct approaches: segmentation-based, anchor-based, keypoint-based, and curve-based strategies. (1) Segmentation-based approaches frame 2D lane detection as a pixel-level classification challenge, generating lane masks[116, 64, 65, 116]. These methods cultivate lane groupings through the exploration of effective semantic features and subsequent post-processing, although they are marked by their high computational costs. (2) Anchor-based approaches are lauded in 2D lane detection for their simplicity and efficiency, typically utilizing linear anchor points to regress positional offsets relative to targets [47, 78, 115]. To circumvent the constraints of linear anchors, [35]employs an intrinsic lane space generating various lane candidates. Heuristically designed row anchors classify row pixels as lanes [109, 50], further evolved into mixed (row and column) anchors in [37] to alleviate positioning errors for side lanes. Moreover, this method significantly enhances inference speed. (3) Keypoint-based approaches offer a more flexible and sparse modeling of lane positions[42, 67, 86], initiating with the estimation of point locations, followed by associating keypoints belonging to the same lane using varying schemes. The primary strategy in [72, 78] is to predict 2D lanes by representing lane structures with predefined keypoints and regressing offsets between sampled and predefined points. Despite promising results, these methods lack the flexibility to adapt to complex lane configurations due to their fixed point design. (4) Curve-based approaches [77, 83, 51]focus on fitting lane lines through various curve equations and specific parameters, transforming 2D lane detection into a curve parameter regression challenge by detecting start and end points alongside curve parameters. Although there have been promising advancements in 2D lane detection, a significant gap remains between the 2D outcomes and the real-world application requirements, notably the precise 3D positioning.
@@
3D lane detection. Due to the inherent lack of depth information in 2D lane detection, projecting these detections into 3D space can result in inaccuracies and reduced robustness[8]. Consequently, numerous researchers have shifted their focus towards lane detection within the 3D domain. Deep learning-based 3D lane detection methodologies primarily bifurcate into CNN-based and Transformer-based approaches, initially constructing dense Bird’s Eye View (BEV) feature maps, followed by the extraction of 3D lane information from these intermediate representations[31].
@@
CNN-based methods predominantly include D-LaneNet [24], which proposes a dual-path architecture utilizing Inverse Perspective Mapping (IPM) to transpose features and detect lanes through vertical anchor regression. 3D-LaneNet+ [16] segments BEV features into non-overlapping units, addressing the limitations of anchor orientation through lateral offsets, angles, and elevation shifts relative to the unit’s center. GenLaneNet [25] pioneers the use of a fictitious top-view coordinate system for better feature alignment and introduces a two-stage framework to decouple lane segmentation and geometric encoding. BEVLaneDet [91] employs a virtual camera to ensure spatial consistency and adapts to more complex scenarios through a keypoint-based representation of 3D lanes. GroupLane [48] innovates within the BEV by introducing a row-based classification method, which accommodates lanes in any direction and interacts with feature information within instance groups.
@@
Transformer-based methods include CLGo [52], presenting a two-stage framework capable of estimating camera posture from the image and lane decoders based on BEV features. PersFormer [12] constructs dense BEV queries using offline camera poses, unifying 2D and 3D lane detection under a Transformer-based framework. STLanes3D [95] predicts 3D lanes using fused BEV features and introduces the 3DLane-IOU loss to couple lateral and height errors. Anchor3DLane [31], a CNN-based method, regresses 3D lanes directly from image features based on 3D anchors, significantly reducing computational overhead. CurveFormer [7] leverages sparse query representations and cross-attention mechanisms within Transformers to effectively regress polynomial coefficients of 3D lanes. LATR [53], building on CurveFormer’s query anchor modeling, constructs lane-aware query generators and dynamic 3D ground position embeddings. CurveFormer++[8] proposes a single-stage Transformer detection method that does not require image feature view transformation and directly infers 3D lane detection results from perspective image features.
@@ I-B Challenges and Motivation
Accurately estimating the 3D position of lane markings demands robust depth perception, particularly in intricate urban settings with fluctuating lighting and weather conditions[43]. Furthermore, real-world data for 3D lane detection exhibits high variability due to factors such as diverse road types, markings, and environmental conditions, making it arduous to train models that generalize well across different scenarios[10]. Processing 3D data for lane detection entails substantial computational resources; this is especially critical in real-time applications where low latency is imperative. Additionally, lane markings may be obscured or degraded by various environmental factors such as occlusions, shadows, rain, or snow, presenting challenges for reliable detection in adverse conditions[113]. Moreover, integrating 3D lane detection into a comprehensive perception system alongside other sensors like cameras, LiDAR, and radar, and processing their combined outputs, poses integration challenges. Unfortunately,the community lacks a consolidated, unified, single point of reference for ascertaining the current level of maturity of camera-based 3D lane detection techniques for autonomous driving.
@@ I-D Review Methodology
The research works discussed in this survey were retrieved using different keywords such as 3D lane detection in autonomous vehicles, vision-based 3D lane detection, and learning-based 3D lane detection. Most of the retrieved papers are directly relevant to the research topic, but there are also some exceptions, such as multi-modal methods [54, 6], and point cloud-based approaches [36], which are less related to the theme of this survey. Additionally, the aforementioned keywords were searched in multiple repositories, including Web of Science and Google Scholar, to ensure the retrieval of relevant content. Inclusion criteria ensured that a paper was recognized by experts in autonomous driving, based on factors such as citation count or the impact of prior work. It is worth mentioning that when consulting the literature, no work on monocular 3D lane detection based on traditional methods was found. This may be because, unlike single-camera 2D lane detection, which only requires identifying pixels belonging to the lane in a 2D image, monocular 3D lane detection requires determining the three-dimensional position information of the lane in the 3D space using the 2D image. Without the assistance of distance measuring sensors such as LiDAR  [^1] https://en.wikipedia.org/wiki/Lidar or prediction through deep learning, this is difficult to achieve.
@@ II MONOCULDAR 3D LANE DETECTION IN AD
With the rapid development of deep learning and autonomous driving, deep learning-based monocular lane detection has received increasing attention from both the industry and academia. In the field of monocular lane detection, early work mainly focused on 2D lane detection. As autonomous driving technology matures, higher requirements are placed on lane detection, with the goal of predicting the three-dimensional information of lane lines from a single image. Therefore, starting from 2018, there have been successive works on monocular 3D lane detection. As shown in Fig. 1, the figure provides a chronological overview of the monocular 3D lane detection algorithms. It can be seen that with the passage of time, more and more works have emerged, indicating that this field is becoming increasingly popular. In this figure, the green arrows represent CNN-based methods, and the orange arrows represent Transformer-based methods.
@@
Among these methods, 3D-LaneNet [24] is a pioneering work in the field of monocular 3D lane detection. 3D-LaneNet introduces a network that can directly predict three-dimensional lane information in road scenes from monocular images. This work is the first to solve the task of three-dimensional lane detection using on-board monocular vision sensor. 3D-LaneNet introduces two novel concepts: intra-network feature map inverse perspective mapping (IPM) and anchor-based lane representation. The intra-network IPM projection facilitates dual representation information flow in both front view and bird’s-eye view. The anchor-based lane output representation supports an end-to-end training approach, which is different from common heuristic methods that equate the problem of detecting three-dimensional lane lines to object detection problems. The overview of 3D-LaneNet is depicted in the Fig. 2. Inspired by works such as FCOS [81] and CenterNet [118], 3D LaneNet+ [16] is an anchor-free 3D lane detection algorithm that can detect 3D lane lines of arbitrary topological structures. The authors of 3D LaneNet+ followed the dual-stream network of 3D LaneNet that separately processes the image view and bird’s-eye view and extended it to support the detection of 3D lane lines with more topological structures. Instead of predicting the entire lane as a whole, 3D LaneNet+ detects small lane segments that lie within the cell and their attributes (position, orientation, height). In addition, the method learns a global embedding for each cell to cluster small lane segments into complete 3D lane information. In [33], the Jiang et al. designed a two-stage 3D lane detection network where each stage is trained separately. The first subnetwork focuses on lane image segmentation, while the second subnetwork focuses on predicting the 3D lane structure based on the segmentation outputs from the first subnetwork. The Efficient Channel Attention (ECA) [88] attention mechanism and the Convolutional Block Attention Module (CBAM) [101] attention mechanism were introduced in each stage, respectively, resulting in improved segmentation performance and accuracy in 3D lane detection. In [25], Guo et al. proposed Gen-LaneNet, a general and scalable 3D lane detection method for detecting 3D lane lines from a single image, as shown in Fig. 3. The authors introduce a novel geometry-guided lane anchor representation and apply specific geometric transformations to directly compute the real 3D lane points from the network output. This anchor design is an intuitive extension of the anchor design in 3D-LaneNet. The method aligns the anchor coordinates with the underlying bird’s-eye-view features, which makes it more robust in handling unfamiliar scenes. Additionally, the paper presents a scalable two-stage framework that allows independent learning of the image segmentation subnetwork and the geometry encoding subnetwork, significantly reducing the number of 3D labels required for training. Furthermore, the paper introduces a highly realistic synthetic image dataset with rich visual variations for the development and evaluation of 3D lane detection methods.
@@
Liu et al. proposed CLGo [52], a two-stage framework for predicting 3D lanes and camera pose from a single image. The first stage focuses on camera pose estimation and introduces an auxiliary 3D lane task and geometry constraints for multi-task learning. The second stage targets the 3D lane task and uses the previously estimated pose to generate top-view images for accurate 3D lane prediction. PersFormer [12] introduces the first transformer-based approach for 3D lane detection and proposes a novel architecture called Perspective Transformer, as shown in Fig. 4. This transformer-based architecture enables spatial feature transformation, allowing for accurate detection of 3D lane lines. Moreover, the proposed framework has the unique capability to simultaneously handle both 2D and 3D lane detection tasks, providing a unified solution. Additionally, the paper presents OpenLane, a large-scale 3D lane detection dataset built upon the influential Waymo Open dataset [75]. OpenLane is the first dataset of its kind to offer high-quality annotations and diverse real-world scenarios, making it a valuable resource for advancing research in the field.

### v2

# reference.bib

@@ Abstract
3D lane detection is essential in autonomous driving as it extracts structural and traffic information from the road in three-dimensional space, aiding self-driving cars in logical, safe, and comfortable path planning and motion control. Given the cost of sensors and the advantages of visual data in color information, 3D lane detection based on monocular vision is an important research direction in the realm of autonomous driving, increasingly gaining attention in both industry and academia. Regrettably, recent advancements in visual perception seem inadequate for the development of fully reliable 3D lane detection algorithms, which also hampers the progress of vision-based fully autonomous vehicles. We believe that there is still considerable room for improvement in 3D lane detection algorithms for autonomous vehicles using visual sensors, and significant enhancements are needed. This review looks back and analyzes the current state of achievements in the field of 3D lane detection research. It covers all current monocular-based 3D lane detection processes, discusses the performance of these cutting-edge algorithms, analyzes the time complexity of various algorithms, and highlights the main achievements and limitations of ongoing research efforts. The survey also includes a comprehensive discussion of available 3D lane detection datasets and the challenges that researchers face but have not yet resolved. Finally, our work outlines future research directions and invites researchers and practitioners to join this exciting field.
@@ I INTRODUCTION
In recent advancements fueled by artificial intelligence, autonomous driving technology has seen rapid evolution, progressively reshaping the paradigms of human transportation. Equipped with an array of sensors, autonomous vehicles emulate human sensory capabilities such as vision and hearing to perceive their surroundings and interpret traffic scenarios for safe navigation[wei2021multi]. Key among these sensors are LiDAR, high-resolution cameras, millimeter-wave radars, and ultrasonic radars[zou2021comparative, ma2022automatic], which facilitate feature extraction and object classification combined with high-precision mapping to discern obstacles and the vehicular traffic landscape[zhao2024curbnet]. Visual sensors, the most extensively utilized in autonomous vehicles, serve as the primary means for environmental perception, encompassing lane detection, traffic light analysis, road sign detection and recognition, vehicle tracking, pedestrian detection, and short-term traffic prediction[chen2021data]. The processing and understanding of visual scenes in autonomous driving, which includes analysis of traffic lights, identification of traffic signs, lane detection, and the detection of nearby pedestrians and vehicles, underpin more robust and safer commands for maneuvers such as steering, overtaking, lane changes, or braking[usman2020speed, ma2024dataset]. This integration of sensory data and environmental understanding seamlessly transitions into the domain of scene understanding in autonomous driving, which is pivotal for advancing vehicular autonomy and ensuring road safety.
@@
Scene understanding represents one of the most challenging aspects within the domain of autonomous driving[ma2023self]. Without the capability for comprehensive scene understanding, navigating an autonomous vehicle safely through traffic lanes can be as daunting as navigating the world blindfolded for humans. Lane detection, in particular, stands out as a pivotal and challenging task within the realm of scene comprehension[xu2023centerlinedet]. Lanes serve as the most common traffic elements on roads, acting as crucial markers that segment the roadways to ensure the safe and efficient passage of vehicles[tang2021review]. Lane detection technology, which automatically identifies road markings, is indispensable; autonomous vehicles lacking this capability could lead to traffic congestion and even severe collisions, thereby compromising passenger safety. Thus, lane detection plays a critical role in the ecosystem of autonomous driving. Unlike typical objects, lane markings occupy only a slender portion of the road scene and are widely distributed, making them uniquely challenging to detect[sun2019accurate]. The task is further complicated by a variety of lane markings, insufficient lighting, obstructions, and interference from similar textures, which are common in many driving scenarios, thus exacerbating the intrinsic challenges associated with lane detection.
@@
Lane detection methods based on monocular vision can be primarily categorized into traditional manual feature approaches and deep learning-based methods[ma2018multiple]. Early endeavors primarily focused on extracting low-level manual features[he2004color, kim2008robust], such as edges and color information. However, these approaches often involve complex feature extraction and post-processing designs, and exhibit limited robustness in dynamically changing scenes. Traditional lane detection algorithms based on manual feature extraction commence by identifying characteristics such as color, texture, edges, orientation, and shape of the lane lines, followed by constructing detection models that approximate the lane markings as straight or higher-order curved lines[zhou2010novel]. Nevertheless, due to the lack of distinguishing features and poor adaptability to dynamic environments, traditional methods based on manual features are typically less reliable and more computationally expensive[wang2004lane]. With the rapid evolution of deep learning, significant strides have been made in the fields of image classification, object detection, and semantic segmentation within computer vision, bringing innovative perspectives to the research of lane detection[wei2022row]. Deep neural networks (DNNs), rooted in deep learning, demonstrate profound capability in feature extraction from image data, with Convolutional Neural Networks (CNNs) being the most extensively applied. CNNs represent a specialized category of DNNs, characterized by multiple convolutional layers and bases, making them particularly suitable for processing structured data like visual images, and providing efficient feature extraction for various subsequent tasks[xu2022rngdet]. In the context of lane detection, this translates to utilizing deep CNNs to extract high-level features in real-time frames, which are then processed by the model to accurately determine the positions of lane lines[huang2023anchor3dlane].
@@ I-A Background and Related Works
Due to advancements in deep learning technologies, researchers have developed numerous strategies that have significantly simplified, expedited, and enhanced the task of lane detection[tabelini2021keep]. Concurrently, as deep learning becomes increasingly widespread and as new concepts continuously emerge[ma20233d], the domain of lane detection has seen a further specialization and refinement in its approaches. Reflecting on the mainstream research trajectories within this domain, camera-based lane detection methods can be principally segmented into two-dimensional (2D) and three-dimensional (3D) lane detection paradigms[chen2022persformer].
@@
2D lane detection. [jin2022eigenlanes, liu2021end] methods aim to accurately delineate the shape and position of lanes within images, primarily employing four distinct approaches: segmentation-based, anchor-based, keypoint-based, and curve-based strategies. (1) Segmentation-based approaches frame 2D lane detection as a pixel-level classification challenge, generating lane masks[zheng2021resa, neven2018towards, pan2018spatial, zheng2021resa]. These methods cultivate lane groupings through the exploration of effective semantic features and subsequent post-processing, although they are marked by their high computational costs. (2) Anchor-based approaches are lauded in 2D lane detection for their simplicity and efficiency, typically utilizing linear anchor points to regress positional offsets relative to targets [li2019line, tabelini2021keep, zheng2022clrnet]. To circumvent the constraints of linear anchors, [jin2022eigenlanes]employs an intrinsic lane space generating various lane candidates. Heuristically designed row anchors classify row pixels as lanes [yoo2020end, liu2021condlanenet], further evolved into mixed (row and column) anchors in [qin2022ultra] to alleviate positioning errors for side lanes. Moreover, this method significantly enhances inference speed. (3) Keypoint-based approaches offer a more flexible and sparse modeling of lane positions[ko2021key, qu2021focus, wang2022keypoint], initiating with the estimation of point locations, followed by associating keypoints belonging to the same lane using varying schemes. The primary strategy in [su2021structure, tabelini2021keep] is to predict 2D lanes by representing lane structures with predefined keypoints and regressing offsets between sampled and predefined points. Despite promising results, these methods lack the flexibility to adapt to complex lane configurations due to their fixed point design. (4) Curve-based approaches [tabelini2004lane, van2019end, liu2021end]focus on fitting lane lines through various curve equations and specific parameters, transforming 2D lane detection into a curve parameter regression challenge by detecting start and end points alongside curve parameters. Although there have been promising advancements in 2D lane detection, a significant gap remains between the 2D outcomes and the real-world application requirements, notably the precise 3D positioning.
@@
3D lane detection. Due to the inherent lack of depth information in 2D lane detection, projecting these detections into 3D space can result in inaccuracies and reduced robustness[bai2024curveformer++]. Consequently, numerous researchers have shifted their focus towards lane detection within the 3D domain. Deep learning-based 3D lane detection methodologies primarily bifurcate into CNN-based and Transformer-based approaches, initially constructing dense Bird’s Eye View (BEV) feature maps, followed by the extraction of 3D lane information from these intermediate representations[huang2023anchor3dlane].
@@
CNN-based methods predominantly include D-LaneNet [garnett20193d], which proposes a dual-path architecture utilizing Inverse Perspective Mapping (IPM) to transpose features and detect lanes through vertical anchor regression. 3D-LaneNet+ [efrat20203d] segments BEV features into non-overlapping units, addressing the limitations of anchor orientation through lateral offsets, angles, and elevation shifts relative to the unit’s center. GenLaneNet [guo2020gen] pioneers the use of a fictitious top-view coordinate system for better feature alignment and introduces a two-stage framework to decouple lane segmentation and geometric encoding. BEVLaneDet [wang2023bev] employs a virtual camera to ensure spatial consistency and adapts to more complex scenarios through a keypoint-based representation of 3D lanes. GroupLane [li2023grouplane] innovates within the BEV by introducing a row-based classification method, which accommodates lanes in any direction and interacts with feature information within instance groups.
@@
Transformer-based methods include CLGo [liu2022learning], presenting a two-stage framework capable of estimating camera posture from the image and lane decoders based on BEV features. PersFormer [chen2022persformer] constructs dense BEV queries using offline camera poses, unifying 2D and 3D lane detection under a Transformer-based framework. STLanes3D [wang2022spatio] predicts 3D lanes using fused BEV features and introduces the 3DLane-IOU loss to couple lateral and height errors. Anchor3DLane [huang2023anchor3dlane], a CNN-based method, regresses 3D lanes directly from image features based on 3D anchors, significantly reducing computational overhead. CurveFormer [bai2022curveformer] leverages sparse query representations and cross-attention mechanisms within Transformers to effectively regress polynomial coefficients of 3D lanes. LATR [luo2023latr], building on CurveFormer’s query anchor modeling, constructs lane-aware query generators and dynamic 3D ground position embeddings. CurveFormer++[bai2024curveformer++] proposes a single-stage Transformer detection method that does not require image feature view transformation and directly infers 3D lane detection results from perspective image features.
@@ I-B Challenges and Motivation
Accurately estimating the 3D position of lane markings based on monocular vision presents numerous challenges. Firstly, real-world data for 3D lane detection is highly variable due to the diversity of weather [kuang2020real], road types, markings, and environmental conditions, which complicates the training of models that can generalize well across different scenarios [campbell2010autonomous, zhang2023perception]. Furthermore, processing 3D data for lane detection requires substantial computational resources, which is particularly critical in autonomous driving applications where low latency is imperative. Reducing the processing time of algorithms without compromising their performance is also a significant challenge in such real-time applications. Recently, the community lacks a consolidated, unified, single point of reference for ascertaining the current level of maturity of camera-based 3D lane detection techniques for autonomous driving.
@@ I-D Review Methodology
The research works discussed in this survey were retrieved using different keywords such as 3D lane detection in autonomous vehicles, vision-based 3D lane detection, and learning-based 3D lane detection. Most of the retrieved papers are directly relevant to the research topic, but there are also some exceptions, such as multi-modal methods [luo2022m, bai2018deep], and point cloud-based approaches [jung2018real], which are less related to the theme of this survey. Additionally, the aforementioned keywords were searched in multiple repositories, including Web of Science and Google Scholar, to ensure the retrieval of relevant content. Inclusion criteria ensured that a paper was recognized by experts in autonomous driving, based on factors such as citation count or the impact of prior work. It is worth mentioning that when consulting the literature, no work on monocular 3D lane detection based on traditional methods was found. This may be because, unlike single-camera 2D lane detection, which only requires identifying pixels belonging to the lane in a 2D image, monocular 3D lane detection requires determining the three-dimensional position information of the lane in the 3D space using the 2D image. Without the assistance of distance measuring sensors such as LiDAR [^1] https://en.wikipedia.org/wiki/Lidar or prediction through deep learning, it is difficult to achieve.
@@ II MONOCULDAR 3D LANE DETECTION IN AD
With the rapid development of deep learning and autonomous driving, deep learning-based monocular lane detection has received increasing attention from both the industry and academia. In the field of monocular lane detection, early work mainly focused on 2D lane detection. As autonomous driving technology matures and the demand for cost reduction increases, higher requirements are placed on lane detection, aiming to predict the three-dimensional information of lane lines from a single image.Therefore, starting from 2018, there have been successive works on monocular 3D lane detection. As shown in Fig. 1, the figure provides a chronological overview of the monocular 3D lane detection algorithms. It can be seen that with the passage of time, more and more works have emerged, indicating that this field is becoming increasingly popular. In this figure, the green arrows represent CNN-based methods, and the orange arrows represent Transformer-based methods.
@@
Among these methods, 3D-LaneNet [garnett20193d] is a pioneering work in the field of monocular 3D lane detection. 3D-LaneNet introduces a network that can directly predict three-dimensional lane information in road scenes from monocular images. This work is the first to solve the task of three-dimensional lane detection using on-board monocular vision sensor. 3D-LaneNet introduces two novel concepts: intra-network feature map inverse perspective mapping (IPM) and anchor-based lane representation. The intra-network IPM projection facilitates dual representation information flow in both front view and bird’s-eye view. The anchor-based lane output representation supports an end-to-end training approach, which is different from common heuristic methods that equate the problem of detecting three-dimensional lane lines to object detection problems. The overview of 3D-LaneNet is depicted in the Fig. 2. Inspired by works such as FCOS [tian2019fcos] and CenterNet [zhou2019objects], 3D LaneNet+ [efrat20203d] is an anchor-free 3D lane detection algorithm that can detect 3D lane lines of arbitrary topological structures. The authors of 3D LaneNet+ followed the dual-stream network of 3D LaneNet that separately processes the image view and bird’s-eye view and extended it to support the detection of 3D lane lines with more topological structures. Instead of predicting the entire lane as a whole, 3D LaneNet+ detects small lane segments that lie within the cell and their attributes (position, orientation, height). In addition, the method learns a global embedding for each cell to cluster small lane segments into complete 3D lane information. In [jiang2022robust], the Jiang et al. designed a two-stage 3D lane detection network where each stage is trained separately. The first subnetwork focuses on lane image segmentation, while the second subnetwork focuses on predicting the 3D lane structure based on the segmentation outputs from the first subnetwork. The Efficient Channel Attention (ECA) [wang2020eca] attention mechanism and the Convolutional Block Attention Module (CBAM) [woo2018cbam] attention mechanism were introduced in each stage, respectively, resulting in improved segmentation performance and accuracy in 3D lane detection. In [guo2020gen], Guo et al. proposed Gen-LaneNet, a general and scalable 3D lane detection method for detecting 3D lane lines from a single image, as shown in Fig. 3. The authors introduced a novel geometry-guided lane anchor representation and applied specific geometric transformations to directly compute the real 3D lane points from the network output. This anchor design extends intuitively from the anchor design in 3D-LaneNet. This method aligns the anchor coordinates with the underlying bird’s-eye-view features, enhancing its robustness in handling unfamiliar scenes. Additionally, this paper presents a scalable two-stage framework enabling independent learning of the image segmentation subnetwork and the geometry encoding subnetwork, thus significantly reducing the number of 3D labels required for training. Furthermore, this paper introduced a highly realistic synthetic image dataset with rich visual variations for developing and evaluating 3D lane detection methods.
@@
Liu et al. proposed CLGo [liu2022learning], a two-stage framework for predicting 3D lanes and camera pose from a single image. Its first stage focuses on camera pose estimation and introduces an auxiliary 3D lane task and geometry constraints for multi-task learning. The second stage targets the 3D lane task and uses the previously estimated pose to generate top-view images for accurate 3D lane prediction. PersFormer [chen2022persformer] introduces the first transformer-based approach for 3D lane detection and proposes a novel architecture called Perspective Transformer, as shown in Fig. 4. This transformer-based architecture enables spatial feature transformation, allowing for accurate detection of 3D lane lines. Moreover, the proposed framework has the unique capability to simultaneously handle both 2D and 3D lane detection tasks, providing a unified solution. Additionally, the paper presents OpenLane, a large-scale 3D lane detection dataset built upon the influential Waymo Open dataset [sun2020scalability]. OpenLane is the first dataset of its kind to offer high-quality annotations and diverse real-world scenarios, making it a valuable resource for advancing research in the field.